# Aeos: Star-by-Star Cosmological Simulations of Early Chemical Enrichment and Galaxy Formation

### v1

@@ 1 Introduction
The first galaxies formed 200–300 million years after the Big Bang, hosting many of the first stars and seeding the creation of every galaxy in the Universe today (Bromm & Yoshida, 2011). The smallest of these early galaxies experienced inefficient star formation for 1–2 Gyr before being quenched by cosmic reionization (Brown et al., 2014), so those that survived until now are composed of ancient stars from 13 Gyr ago. Stars from these ultra-faint dwarf galaxies ($L<10^{5}L_{\odot}$) are therefore relics from the era of the first stars and galaxies, preserving signatures of early chemical enrichment (Bovill & Ricotti, 2009; Simon, 2019).
@@
Chemical evolution models that rely on simple parametric approaches and assume homogeneous mixing are capable of reproducing mean trends but incapable of modeling scatter (e.g., Andrews et al., 2017; Côté & Ritter, 2018). When chemical abundance observations of dwarf galaxies were sparse, these models were sufficient to explain observed trends and differences between dwarf galaxies as a function of luminosity (e.g., Tolstoy et al., 2009; Kirby et al., 2011). Recent observations, however, have revealed abundance scatter in many elements both within and between dwarf galaxies (e.g., Hill et al., 2019; Ji et al., 2020; Mead et al., 2024). Such variations from mean trends imply the presence of complex galaxy formation processes like source- and time-dependent metal mixing, hierarchical galaxy merging, and bursty star formation (e.g., Emerick et al., 2020).
@@
More recently, hydrodynamic galaxy simulations have begun to reach several parsec resolution and include star particles representing individual stars (e.g., Emerick et al., 2019; Hirai et al., 2021; Gutcke et al., 2021; Hislop et al., 2022; Andersson et al., 2023; Lahén et al., 2020; Steinwandel et al., 2023; Calura et al., 2022; Deng et al., 2024). This level of detail allows simulations to capture variations in when, where, and how individual stars inject energy and metals. Such detail in a simulation allows investigation into the origins of the full chemical abundance distributions observed in dwarf galaxies beyond just the mean elemental trends. Revaz et al. (2016) showed that below mass resolution of 103 M, stellar populations are not representative of the full IMF. This leads to oversimplifications in enrichment and the inability to properly model abundance scatter. This was similarly argued by Smith (2021) for mass resolution below 500 M, finding that stochastic variations in stellar populations become important when modeling radiation and galactic wind driving. In the smallest galaxies, this becomes especially important, as each individual feedback event affects the evolution of the galaxy. To understand the origin of these distributions and interpret both present and future chemical abundance data, simulations need to effectively combine extremely high-resolution star particles and gas mixing with detailed element yields and metal tracing techniques.
@@ 2.1.1 Hydrodynamics
Unlike in E19, we do not use an artificial pressure floor, relying instead on the efficient conversion of dense gas into stars and dispersion during subsequent feedback events to prevent the accumulation of numerically unstable dense gas.
@@ 2.1.2 Radiative Cooling and Chemistry
New in the Aeos simulations is: 1) the additional contribution of the UV background to the FUV band, 2) the use in the photoelectric heating rate of a constant efficiency parameter $\epsilon=0.05$ instead of one that depends on local gas density, 3) the effects that both FUV and infrared (IR) radiation have on H2 and H- reaction rates, and 4) a UV background that has been extended to high redshift in all bands by adopting rates assuming a blackbody radiation spectrum at 3×$\times 10^{4}$ K that turns on at $z=50$ and is scaled to be continuous with Haardt & Madau (2012) at $z=10$. In addition, we use an updated Lyman-Werner (LW) background model from both E19 and Wise et al. (2012a), adopting the rates at high redshift from Qin et al. (2020).
@@ 2.1.3 Feedback Injection
Similarly, stellar winds are deposited in the same 2 pc radius region. The mass loss rates for both AGB and massive star winds are adopted from stellar evolution models, but are assumed to have fixed velocities over their lifetime.
@@ 2.3.3 Pop II Stellar Feedback
However, this table does not have complete coverage over all possible stellar properties encountered in these simulations, particularly for stars below about 15 M and very massive stars with sub-solar metallicity. For stars off of the grid, we adopt a blackbody spectrum with rates scaled to be continuous with the OSTAR2002 grid (see Appendix B of E19). Ionizing photon energies are taken to be the average ionizing photon energy for the corresponding blackbody spectrum of each star. Stellar wind velocities are fixed to 20 km s-1 for AGB stars ($M_{*}<8$ M), and 100 km s-1 (our wind velocity ceiling, see Section 2.3) for massive stars ($M_{*}>8$ M). CCSN occur for stars between 8–25 M with an energy of 1051 erg, and we assume stars more massive than 25 M directly collapse into black holes with no mass or energy feedback.
@@ 2.4.2 Pop II Yields
We adopted our particular set of elemental yields based on comparing the results of a one-zone galactic chemical evolution model as applied to a Milky Way mass galaxy. While these yields generally produced reasonable agreement in this model as compared to observations in [X/Fe] vs. [Fe/H] space, Mg is noticeably under-produced in [Mg/Fe] at all [Fe/H], while other $\alpha$-elements tend to agree well with observed Milky Way chemical abundances. Given the importance of accurate observational comparisons in our simulations, we applied a uniform factor of 2.2 to the Mg yields from all massive stars, which brings the model into agreement with observed Mg abundances.
@@ 2.4.3 SNIa Yields
As discussed in 2.3.3, we use a combined DTD from four different sources of SNIa, each with potentially unique abundance signatures. However, given the uncertainty in yields from each of these sources, we opt instead to make the assumption that there is one yield pattern for each source, which allows us to post-process the abundance patterns from each channel separately. Live in the simulation, we assume a single abundance pattern for all SNIa from Thielemann et al. (1986) and track the contribution of each SNIa type to the total metallicity as a separate passive scalar tracer field. Given this, and knowing the total number of each SNIa type that has occurred in the simulation, one can arbitrarily re-scale the abundance patterns for each SNIa type. Doing so implicitly assumes that the yields for each SNIa do not affect the dynamical evolution of our galaxies (which, in reality, it may, for example by influencing cooling through the Fe atomic line cooling), since we do not account for local cooling variations due to individual elemental abundances in our simulations.
@@ 3 Overview of Fiducial Simulation
Of the 91 star-forming halos, the vast majority are tiny. 60 of the halos contain stellar masses of $M_{*}\leq 100$ M, while only one galaxy has $M_{*}>10^{6}$ M. 17 of the galaxies have begun forming Pop II stars by $z\sim 14.5$. For each of these galaxies, the distribution of Pop II vs. Pop III stellar mass is shown in Figure 3.
@@ 3.1 External Enrichment and Systems of Galaxies
Due to the shallow potentials of their small halos, the Aeos galaxies struggle to retain their gas and metals (Mead et al., in prep). Every SN explosion, in the smallest halos, blows out the galaxy’s gas, causing neighboring galaxies to freely share metals and gas between themselves. This leads to behavior like the external enrichment shown in Figure 4.
@@
Figure 5: (Left) Around 160 Myr after the Big Bang, three small galaxies merge in the simulation. Each progenitor enters with its own stars, with colors labeled in the right panel, and the fully merged galaxy continues to form stars. (Right) We show the birth time and metallicity of each star in this system; the galaxies formed stars at slightly different metallicities prior to merging. Each star is colored by its galaxy of origin. We investigate more chemical abundances in Figure 6 to show that structure in chemical abundance space can correspond to the origins of the stars.
@@ 4.1 Structure in Chemical Abundance Space Corresponding to Galactic Origins of Stars
As a proof of concept, consider the merger of three galaxies shown in Figure 5. In simulation, three early galaxies merge at $z\sim 15$. These galaxies contain only 6000 M, 2000 M, and 15000 M of stellar mass respectively before merging, and after merging experience bursty star formation until the merged halo has a stellar mass of $10^{7}$ M at the end of the simulation (Halo 1 at $z=14.5$). This is a clustered region of the simulation.
Figure 5 shows the individual star particles of these galaxies, colored according to their progenitor galaxy. As the galaxies merge, star formation increases and the merged galaxy continues to form stars (shown in orange). Here, “fully merged” means that the star-forming regions are indistinguishable from one another at the physical scale of 0.1 kpc. The projections and stellar birth times shown in Figure 5 can only ever be known in a simulation, never in any observed data.
The stellar chemical abundances shown in Figure 6, however, could be measured in observed stars with spectroscopy. This figure shows the [Fe/H], [C/Fe], [N/Fe], [O/Fe], and [Sr/Fe] chemical abundances of the low-mass Pop II stars that would survive to the present day. To mimic the inherent limitations of observations, the abundances are binned according to typical observational abundance uncertainties for stars in the Milky Way halo – here, we use the typical uncertainties of $\sim$0.1 dex for Fe, $\sim$0.2 dex for Sr, $\sim$0.1-0.2 dex for O (higher at more metal-poor), and $\sim$0.3 dex for N and C. The size of the dot in each bin is proportional to the log number of stars in that bin. Once again, we color the stars according to their progenitor galaxy.
At this point in the simulation in this system, the most important nucleosynthetic sources for the elements in these stars are CCSN (both Pop II and Pop III) with a minimal contribution from AGB winds. In the last $\sim 15$ Myr of the simulation, Pop II CCSN take over as the dominant source of metals in these galaxies. For the elements shown in Figure 6, carbon has the greatest AGB contribution, but the mass of carbon from AGB stars is still almost four orders of magnitude lower than the mass from CCSN (see Section 4.2). The spread in strontium relative to iron is due to strontium’s more significant production in Pop II AGB winds and CCSN compared to Pop III CCSN (see Figure A.1), exacerbated by the overall extremely low amounts of strontium in the gas ($\sim 10^{-4}$ M of strontium in the entire system). Note that strontium is primarily a tracer of AGB winds but forms in small amounts in CCSN due to weak s-process, and the metal content of Pop II stars provide seed nuclei that Pop III stars lack. As time continues to pass, winds and SNIa will become significant and contribute to more scatter, even within individual progenitor galaxies (such as the spread in strontium seen in Figure 6).
@@ 4.2 CEMP Signature in First Generation Pop II Stars
The plot in Figure 7 provides a detailed view of the CEMP fraction across a range of metallicities for all the low-mass stars in the simulation. It shows the fraction of stars with [C/Fe] > 1.0 (blue) and [C/Fe] > 0.7 (green) as a function of [Fe/H]. At the lowest metallicities ([Fe/H] < -5), nearly all low-mass stars in the simulation exhibit a CEMP signature, in alignment with the idea that carbon-enhancement is common in metal-poor environments due to enrichment from the first stars (Pop III) through supernovae or winds. As metallicity increases beyond [Fe/H] $\sim$ -4, the fraction of CEMP stars drops significantly, reflecting the shift from environments dominated by Pop III chemical signatures to those where normal Pop II star formation has begun occurring.
In Figure 7, we also compare to the observed frequencies of CEMP stars in the Milky Way (Placco et al., 2014). At the lowest metallicities, we under-predict CEMP stars with [C/Fe] > 1.0 when compared to Milky Way observations. This could indicate the need for a Pop III IMF that favors higher-mass Pop III, because higher mass CCSNe produce more carbon-rich ejecta. Our Pop III IMF has a characteristic mass of $M_{\text{char}}=10$ M, which is on the lower end of expectations for Pop III IMF. It could also indicate too-efficient mixing in the simulation volume, which could dilute the carbon enhancement in localized star-forming regions. We note, however, that the low-metallicity end of observations is uncertain due to limited data and the simulated abundances generally agree with observed rates of CEMP stars with [C/Fe] > 0.7.

### v2

@@ 1 Introduction
The first galaxies formed 200–300 million years after the Big Bang, hosting many of the first stars and seeding the creation of every galaxy in the Universe today (Bromm & Yoshida, 2011). The smallest of these early galaxies formed stars for 1–2 Gyr before being quenched by cosmic reionization and stellar feedback (Brown et al., 2014), so those that survived until now are composed of ancient stars from 13 Gyr ago. Stars from these ultra-faint dwarf galaxies ($L<10^{5}L_{\odot}$) are therefore relics from the era of the first stars and galaxies, preserving signatures of early chemical enrichment (Bovill & Ricotti, 2009; Simon, 2019).
@@
Chemical evolution models that rely on simple parametric approaches and assume homogeneous mixing are capable of reproducing mean trends but incapable of modeling scatter (e.g., Andrews et al., 2017; Côté & Ritter, 2018). When chemical abundance observations of dwarf galaxies were sparse, these models were sufficient to explain observed trends and differences between dwarf galaxies as a function of luminosity (e.g., Tolstoy et al., 2009; Kirby et al., 2011). Recent observations, however, have revealed abundance scatter in many elements both within and between dwarf galaxies (e.g., Hill et al., 2019; Ji et al., 2020; Mead et al., 2024a). Such variations from mean trends imply the presence of complex galaxy formation processes like source- and time-dependent metal mixing, hierarchical galaxy merging, and bursty star formation (e.g., Emerick et al., 2020).
@@
More recently, hydrodynamic galaxy simulations have begun to reach parsec resolution and include star particles representing individual stars (e.g., Emerick et al., 2019; Hirai et al., 2021; Gutcke et al., 2021; Hislop et al., 2022; Andersson et al., 2023; Lahén et al., 2020; Steinwandel et al., 2023; Calura et al., 2022; Deng et al., 2024). This level of detail allows simulations to capture variations in when, where, and how individual stars inject energy and metals. Such detail in a simulation allows investigation into the origins of the full chemical abundance distributions observed in dwarf galaxies beyond just the mean elemental trends. Revaz et al. (2016) showed that below mass resolution of 103 M, stellar populations are not representative of the full IMF. This leads to oversimplifications in enrichment and the inability to properly model abundance scatter. This was similarly argued by Smith (2021) for mass resolution below 500 M, finding that stochastic variations in stellar populations become important when modeling radiation and galactic wind driving. In the smallest galaxies, this becomes especially important, as each individual feedback event affects the evolution of the galaxy. To understand the origin of these distributions and interpret both present and future chemical abundance data, simulations need to effectively combine extremely high-resolution star particles and gas mixing with detailed element yields and metal tracing techniques.
@@ 2.1.1 Hydrodynamics
Unlike in E19, we do not use an artificial pressure floor. Instead, our simulations operate at sufficiently high resolution, allowing us to accurately resolve dense gas regions. This high resolution enables the efficient conversion of dense gas into stars and ensures that stellar feedback processes, such as supernova explosions and stellar winds, effectively disperse the remaining gas.
@@ 2.1.2 Radiative Cooling and Chemistry
New in the Aeos simulations is: 1) the additional contribution of the UV background to the FUV band, 2) the use in the photoelectric heating rate of a constant efficiency parameter instead of one that depends on local gas density, 3) the effects that both FUV and infrared (IR) radiation have on H2 and H- reaction rates, and 4) a UV background that has been extended to high redshift in all bands by adopting rates assuming a blackbody radiation spectrum at 3×$\times 10^{4}$ K that turns on at $z=50$ and is scaled to be continuous with Haardt & Madau (2012) at $z=10$. The UV background is included to account for the radiation from stars beyond the edges of the simulated volume. In addition, we use an updated Lyman-Werner (LW) background model from both E19 and Wise et al. (2012a), adopting the rates at high redshift from Qin et al. (2020).
For the photoelectric heating rate, we estimate the efficiency as $\epsilon=0.05$. Calculating the photoelectric heating efficiency requires a determination of the electron number density $n_{e}$, which is challenging in dense and neutral regions. In these environments, $n_{e}$ is largely influenced by the ionization of carbon, dust grains, and polycyclic aromatic hydrocarbons. Our existing chemical network only accounts for electrons contributed by hydrogen, helium, and molecular hydrogen. To address this gap, we considered a power-law relationship for the heating efficiency $\epsilon$ based on the hydrogen number density $n_{H}$, as derived from the Wolfire et al. (2003) model of $\Gamma_{\text{Pe}}$ in the solar neighborhood, as also done in E19: $\epsilon=0.0148\times n_{H}^{0.235}$. For densities ranging from 1 to 10000 cm-3, this ranges from $\epsilon$ = 0.015 to 0.12, with typical values around 0.05. We thus take 0.05 as our efficiency.
@@ 2.1.3 Feedback Injection
Similarly, stellar winds are deposited in the same 2 pc radius region. The mass loss rates for both AGB and massive star winds are adopted from stellar evolution models, but are assumed to have fixed velocities and constant loss rates over their AGB phase or lifetime.
@@ 2.3.3 Pop II Stellar Feedback
However, this table does not have complete coverage over all possible stellar properties encountered in these simulations, particularly for stars below about 15 M and very massive stars with sub-solar metallicity. For stars off of the grid, we adopt a blackbody spectrum with rates scaled to be continuous with the OSTAR2002 grid (see Appendix B of E19). Ionizing photon energies are taken to be the average ionizing photon energy for the corresponding blackbody spectrum of each star. Stellar wind velocities are fixed to 20 km s-1 for AGB stars ($M_{*}<8$ M), and 100 km s-1 (our wind velocity ceiling, see Section 2.3) for massive stars ($M_{*}>8$ M). CCSN occur for stars between 8–25 M with an energy of 1051 erg, and we assume stars more massive than 25 M directly collapse into black holes with no mass or energy feedback (Limongi & Chieffi, 2018).
@@ 2.4.2 Pop II Yields
We adopted our particular set of elemental yields based on comparing the results of a one-zone galactic chemical evolution model as applied to a Milky Way mass galaxy. While these yields generally produced reasonable agreement in this model as compared to observations in [X/Fe] vs. [Fe/H] space, Mg is noticeably under-produced in [Mg/Fe] at all [Fe/H], while other $\alpha$-elements tend to agree well with observed Milky Way chemical abundances. Given the importance of accurate observational comparisons in our simulations, we applied a uniform factor of 2.2 to the Mg yields from all massive stars, which brings the model into agreement with observed Mg abundances. This is discussed in Appendix A.
@@ 2.4.3 SNIa Yields
As discussed in 2.3.3, we use a combined DTD from four different sources of SNIa, each with potentially unique abundance signatures. However, given the uncertainty in yields from each of these sources, we opt instead to make the assumption that there is one yield pattern for each source, which allows us to post-process the abundance patterns from each channel separately. Live in the simulation, we assume a single abundance pattern for all SNIa from Thielemann et al. (1986) and track the contribution of each SNIa type to the total metallicity as a separate passive scalar tracer field. Given this, and knowing the total number of each SNIa type that has occurred in the simulation, one can arbitrarily rescale the abundance patterns for each SNIa type. Doing so implicitly assumes that the yields for each SNIa do not affect the dynamical evolution of our galaxies (which, in reality, it may, for example by influencing cooling through the Fe atomic line cooling), since we do not account for local cooling variations due to individual elemental abundances in our simulations.
@@ 3 Overview of Fiducial Simulation
Some galaxies in our simulations exhibit unexpectedly high stellar-to-halo mass ratios (such as $M_{*}=10^{5}\,M_{\odot}$ for $M_{\text{halo}}=10^{6}\,M_{\odot}$). Abundance matching relations are very poorly calibrated for low-mass halos, but they predict closer to $M_{*}=10^{2}-10^{3}\,M_{\odot}$ or even lower for a $10^{6}\,M_{\odot}$ halo (e.g., Behroozi et al., 2013). Most of our galaxies are indeed low in stellar mass, and we have significant scatter in stellar mass, which is expected with lower halo masses (Garrison-Kimmel et al., 2017). For the massive outliers, their high stellar-to-halo mass ratios may be due in part to environmental factors. Two of the outlier galaxies are the externally enriched galaxies near the central galaxy (see Section 3.1), and their total halo mass is affected by their orbit around a more massive galaxy, allowing for dark matter stripping (a lesser version of the stripping processes described by Moreno et al. (2022)).
The unexpectedly high stellar-to-halo mass ratios may also be due to our star-by-star IMF sampling approach. Recently, Jeon & Ko (2024) found that star-by-star sampling leads to higher stellar masses in low-mass halos (e.g., $10^{8}\,M_{\odot}$) compared to traditional single stellar population (SSP) particles, which inject feedback more burstily and suppress star formation more effectively. Jeon & Ko (2024) demonstrated that the SSP method results in burstier, stronger feedback, while our star-by-star method, with its weaker feedback, allows for higher stellar masses. They also argue that star-by-star sampling more accurately matches observations and accounts for the high scatter in the $M_{\text{halo}}$-$M_{*}$ relation at low halo masses. Similarly, Andersson et al. (2024) reports a systematic increase in stellar mass for a given halo mass compared to classical galaxy models (see Figure 1 of Andersson et al. (2024)). In our other recent Aeos paper (Mead et al., 2024b), we observe that a single SN disrupts star formation in most galaxies, but with this effect diminishes around $10^{7}\,M_{\odot}$ where our outliers lie.
Of the 91 star-forming halos, the vast majority are tiny. 60 of the halos contain stellar masses of $M_{*}\leq 100$ M, while only one galaxy has $M_{*}>10^{6}$ M. 17 of the galaxies have begun forming Pop II stars by $z\sim 14.5$. For each of these galaxies, the distribution of Pop II vs. Pop III stellar mass is shown in Figure 3. Galaxies typically undergo multiple episodes of Pop III star formation before transitioning to Pop II stars. The extent of Pop III star formation required for this transition is strongly influenced by the halo mass, which affects the galaxy’s ability to retain metals, as well as the star formation activity of neighboring galaxies. For example, in Halo 1 of Figure 3, approximately 1,000 solar masses of Pop III stars formed over an 80 Myr period before the onset of Pop II star formation. Other galaxies exhibited a different amounts of Pop III star formation, ranging from none or only a few hundred solar masses – particularly in cases of external enrichment, more massive halos, or proximity to other star-forming galaxies – to up to about 1,000 solar masses in galaxies that began star formation earlier than their surroundings or were relatively isolated. This variation highlights the significant role that both intrinsic halo properties and environmental factors play in determining the transition from Pop III to Pop II star formation.
@@ 3.1 External Enrichment and Systems of Galaxies
Due to the shallow potentials of their small halos, the Aeos galaxies struggle to retain their gas and metals (Mead et al., 2024b). Every SN explosion, in the smallest halos, blows out the galaxy’s gas, causing neighboring galaxies to freely share metals and gas between themselves. This leads to behavior like the external enrichment shown in Figure 4.
@@
We also look at how the metallicities differ for stars formed via external enrichment. When looking at first-generation Pop II stars in the galaxies shown in Figure 4, the metallicities differ between the externally enriched halos and the central halo, but can be both higher metallicity or lower metallicity. The first Pop II stars in Halo 1 form at about [Fe/H] = $-5$, while the first stars in Halo 3 form at about [Fe/H] = $-4.5$. When SNe blow out the metals from Halo 1, it actually results in Halo 3 having a higher metal density. On the other hand, Halo 5 is about twice as far away ($\sim 3$ kpc), and when it starts forming stars later at 260 Myr its stars have [Fe/H] = $-5.3$. In this case, the metals became diffused during the transfer.
@@ 4.1 Structure in Chemical Abundance Space Corresponding to Galactic Origins of Stars
As a proof of concept, consider the merger of three galaxies shown in Figure 5. In simulation, three early galaxies merge at $z\sim 15$. At the time of merger, these galaxies contain only 6000 M, 2000 M, and 15000 M of stellar mass respectively before merging, and after merging experience bursty star formation until the merged halo has a stellar mass of $10^{7}$ M at the end of the simulation (Halo 1 at $z=14.5$). This is a clustered region of the simulation.
Figure 5: (Left) Around 160 Myr after the Big Bang, three small galaxies merge in the simulation. Each progenitor enters with its own stars, with colors labeled in the right panel, and the merged galaxy continues to form stars. (Right) We show the birth time and metallicity of each star in this system; the galaxies formed stars at slightly different metallicities prior to merging. Each star is colored by its galaxy of origin. We investigate more chemical abundances in Figure 6 to show that structure in chemical abundance space can correspond to the origins of the stars.
@@
Figure 5 shows the individual star particles of these galaxies, colored according to their progenitor galaxy. As the galaxies merge, star formation increases and the merged galaxy continues to form stars (shown in orange). Here, “merged” means that the star-forming regions are indistinguishable from one another at the physical scale of 0.1 kpc, but note that the regions are still oscillating and the halos could still move out and back into Halo 1 before becoming permanently indistinguishable. The projections and stellar birth times shown in Figure 5 can only ever be known in a simulation, never in any observed data.
The stellar chemical abundances shown in Figure 6, however, could be measured in observed stars with spectroscopy. This figure shows the [Fe/H], [C/Fe], [N/Fe], [O/Fe], and [Sr/Fe] chemical abundances of the low-mass Pop II stars that would survive to the present day. To mimic the inherent limitations of observations, the abundances are binned according to typical observational abundance uncertainties for stars in the Milky Way halo – here, we use the typical uncertainties of $\sim$0.1 dex for Fe, $\sim$0.2 dex for Sr, $\sim$0.1-0.2 dex for O (higher at more metal-poor), and $\sim$0.3 dex for N and C. These uncertainties are based on uncertainties of Milky Way stars from JINAbase (Abohalima & Frebel, 2018). The size of the dot in each bin is proportional to the log number of stars in that bin. Once again, we color the stars according to their progenitor galaxy.
At this point in the simulation in this system, the most important nucleosynthetic sources for the elements in these stars are CCSN (both Pop II and Pop III) with a minimal contribution from AGB winds. In the last $\sim 15$ Myr of the simulation, Pop II CCSN take over as the dominant source of metals in these galaxies. For the elements shown in Figure 6, carbon has the greatest AGB contribution, but the mass of carbon from AGB stars is still almost four orders of magnitude lower than the mass from CCSN (see Section 4.2). The spread in strontium relative to iron is due to strontium’s more significant production in Pop II AGB winds and CCSN compared to Pop III CCSN (see Figure A.1), exacerbated by the overall extremely low amounts of strontium in the gas ($\sim 10^{-4}$ M of strontium in the entire system). Note that strontium is primarily a tracer of AGB winds but forms in small amounts in CCSN due to weak s-process, and the metal content of Pop II stars provide seed nuclei that Pop III stars lack. As time continues to pass, winds and SNIa will become significant and contribute to more scatter, even within individual progenitor galaxies (such as the spread in strontium seen in Figure 6). There is also a spread seen in [N/Fe] in the different progenitor galaxies, particularly Progenitor 2. This is due to yield differences from different-mass CCSNe, producing an intrinsic source of scatter.
@@ 4.2 CEMP Signature in First Generation Pop II Stars
The plot in Figure 7 provides a detailed view of the CEMP fraction across a range of metallicities for all the low-mass stars in the simulation. It shows the fraction of stars with [C/Fe] > 1.0 (blue) and [C/Fe] > 0.7 (green) as a function of [Fe/H]. At the lowest metallicities ([Fe/H] < -5), nearly all low-mass stars in the simulation exhibit a CEMP signature, in alignment with the idea that carbon-enhancement is common in metal-poor environments due to enrichment from the first stars (Pop III) through faint supernovae or winds. As metallicity increases beyond [Fe/H] $\sim$ -4, the fraction of CEMP stars drops significantly, reflecting the shift from environments dominated by Pop III chemical signatures to those where normal Pop II star formation has begun occurring.
In Figure 7, we also compare to the observed frequencies of CEMP stars in the Milky Way (Placco et al., 2014). At the lowest metallicities, we under-predict CEMP stars with [C/Fe] > 1.0 when compared to Milky Way observations. This could indicate the need for a Pop III IMF that favors higher-mass Pop III, because higher mass CCSNe produce ejecta with a higher ratio of carbon to iron (see Figure A.2 and Heger & Woosley, 2010). Our Pop III IMF has a characteristic mass of $M_{\text{char}}=10$ M, which is on the lower end of expectations for Pop III IMF. It could also indicate too-efficient mixing in the simulation volume, which could dilute the carbon enhancement in localized star-forming regions. We note, however, that the low-metallicity end of observations is uncertain due to limited data and the simulated abundances generally agree with observed rates of CEMP stars with [C/Fe] > 0.7.